### v1

# LLM-Crowdsourced: A Benchmark-Free Paradigm for Mutual Evaluation of Large Language Models

Abstract
Although large language models (LLMs) demonstrate remarkable capabilities across various tasks, evaluating their capabilities remains a challenging task. Existing evaluation methods suffer from issues such as data contamination, black-box operation, and subjective preference. These issues make it difficult to evaluate the LLMs’ true capabilities comprehensively. To tackle these challenges, we propose a novel benchmark-free evaluation paradigm, LLM-Crowdsourced. It utilizes LLMs to generate questions, answer independently, and evaluate mutually. This method integrates four key evaluation criteria: dynamic, transparent, objective, and professional, which existing evaluation methods cannot satisfy simultaneously. Experiments on eight mainstream LLMs across mathematics and programming verify the advantages of our method in distinguishing LLM performance. Furthermore, our study reveals several novel findings that are difficult for traditional methods to detect, including but not limited to: (1) Gemini demonstrates the highest original and professional question-design capabilities among others; (2) Some LLMs exhibit “memorization-based answering” by misrecognizing questions as familiar ones with a similar structure; (3) LLM evaluation results demonstrate high consistency (robustness).
Figure 1: Evaluation criteria and methodologies of existing approaches and our method.
Codehttps://github.com/Qhovo1/LLM-Crowdsourced
Introduction
In recent years, large language models (LLMs) have demonstrated remarkable general capabilities in various scenarios (Achiam et al. 2023; Team et al. 2023; Touvron et al. 2023; Ge et al. 2023), including scientific research assistance (Boyko et al. 2023), code generation (Nejjar et al. 2025), and educational question answering (Kasneci et al. 2023), and have become a central pillar of artificial intelligence advancement (Vaswani et al. 2017). As both the scale of these LLMs and the scope of their applications expand, the differences in the capabilities of various LLMs across different tasks have become increasingly apparent (Chang et al. 2024). While they excel at tasks such as mathematical reasoning (Ahn et al. 2024; Azerbayev et al. 2023; Luo et al. 2023; Lu et al. 2023; Chern et al. 2023; Yu et al. 2023) and code generation (Xu et al. 2022), their abilities are still uneven across different domains. Some LLMs excel at logical reasoning, while others are more proficient in engineering implementation or language understanding. These differences not only impact the selection of LLMs for practical applications but also raise the bar for comparison and development of LLMs. Consequently, there is an urgent need for a systematic, objective, and discriminative evaluation paradigm to reflect the performance of LLMs across multidimensional tasks accurately.
Existing evaluation methods for LLMs mainly include white-box evaluation, black-box evaluation, and human-crowdsourced evaluation. White-box evaluation relies on publicly available and transparent static datasets, offering strong reproducibility and transparency. However, as LLM performance continues to improve, many benchmark datasets have become saturated (Kiela et al. 2021; Vendrow et al. 2025), with multiple LLMs achieving nearly perfect scores, which weakens the discriminative power and utility of these evaluations (Etzine et al. 2025; Millière and Buckner 2024). Furthermore, due to long-term public access to these datasets, LLMs may have encountered related samples during their training phase, leading to data contamination (Cheng, Chang, and Wu 2025; Xu et al. 2024; Deng et al. 2023) and impacting the credibility of the evaluation results (Yang et al. 2023; Sainz et al. 2023). Black-box evaluation mitigates the risk of data contamination by concealing the test set; however, the evaluation process suffers from issues related to black-box operation, lacking credibility and verifiability (Casper et al. 2024). Human-crowdsourced evaluation establishes rankings through user preference voting (Gudibande et al. 2023; Chiang et al. 2024). While this approach avoids the information leakage risks associated with fixed question banks, it heavily depends on users’ subjective preference, making it challenging to ensure the professionalism of the questions and the verifiability of the evaluation process (Connor Chen and Tianle Li 2025; Li, Angelopoulos, and Chiang 2024).
This paper proposes a novel benchmark-free evaluation paradigm based on LLM-Crowdsourced. The core concept of this approach is to have multiple LLMs take turns generating question, answer independently, and evaluate mutually, with final capability rankings computed by aggregating scores across rounds.
In terms of design, this mechanism offers several advantages: questions are dynamically generated by the LLMs, which avoids overlap between the question bank and training data, and significantly mitigates data contamination. The entire evaluation process is fully public and supports review, thereby enhancing transparency. Furthermore, using multi-LLM evaluation in place of a single evaluation LLM or human evaluation effectively mitigates the influence of subjective preference. Both question generation and evaluation are performed independently by LLMs, which, compared to human-crowdsourced, possess more specialized background knowledge. Thus, they can provide more professional and reliable evaluation results.
To verify our proposed LLM-Crowdsourced evaluation paradigm, we systematically evaluated eight mainstream LLMs (such as gpt-4.1, gemini-2.5-pro, deepseek-r1, etc.) in two classic domains: mathematics and programming. The results show that our method not only effectively distinguishes differences among LLMs in logical reasoning, engineering capability, and generalization ability, but also reveals multiple novel findings that are difficult to detect via traditional methods. For example, some LLMs demonstrate a strong ability to generate mathematical questions, being able to construct original questions with theoretical depth and challenge, reflecting a high level of professionalism. We also observed “memorization-based answering” behaviors in some LLMs during the experiments. In addition, the evaluation results among LLMs demonstrate a high consistency, validating the robustness and objectivity of our proposed “decentralized multi-LLM evaluation” mechanism.
Our contributions are summarized as follows:
  1. 1.
    New Evaluation Paradigm: We propose an evaluation paradigm based on LLM-Crowdsourced, which integrates four key evaluation criteria: dynamic, transparent, objective, and professional, effectively overcoming the limitations of current evaluation methods in terms of data contamination, lack of transparency, and insufficient objectivity and professionalism.
  2. 2.
    Novel Findings: We systematically experiment with eight mainstream LLMs in the two classic tasks of mathematics and programming, validating the effectiveness of our proposed evaluation paradigm in differentiating LLMs’ logical reasoning, engineering implementation, and generalization abilities. Meanwhile, the experiments reveal several phenomena that are difficult for traditional methods to detect, including some LLMs demonstrating highly original and professional question-generation capabilities, the existence of “memorization-based answering” behavior, and a high consistency in evaluation results among LLMs.
  3. 3.
    Open-sourced Project: We have publicly released the complete evaluation framework and experimental code to support the community in testing newly released LLMs and to advance the standardization and progress of LLM evaluation methods.
Related Work and Motivations
The evaluation of LLMs is intended to objectively assess LLM performance and generate rankings through multiple evaluation and analysis methods. Evaluation results not only demonstrate the strengths and weaknesses of the LLMs but also provide clear references for researchers, developers, and users. Currently, evaluation methods for LLMs can be categorized into three main types: white-box evaluation, black-box evaluation, and human-crowdsourced evaluation.
White-Box
White-box evaluation is currently the most widely used method for evaluating LLMs. It relies on publicly available and transparent professional datasets with objective standard answers to evaluate LLM performance on specific tasks. Since the test content, evaluation criteria, and procedures are all open and transparent, this method exhibits excellent transparency, low cost, and reproducibility. Therefore, it is widely adopted for the horizontal comparison and research analysis of LLM capabilities. Stanford University first introduced a systematic white-box evaluation framework, HELM (Liang et al. 2023), which selected 16 core tasks and standard datasets to evaluate 30 mainstream LLMs, thus setting a standard and providing a precedent for follow-up research. Current white-box evaluation covers a wide range of task domains, including mathematical reasoning (e.g., Math (Hendrycks et al. 2021), GSM-8K (Cobbe et al. 2021)), code generation (e.g., HumanEval (Chen et al. 2021), LiveCodeBench (Jain et al. 2024)), multidisciplinary question answering (e.g., MMLU (Hendrycks et al. 2020), AGIEval (Zhong et al. 2023)), and commonsense inference (e.g., HellaSwag (Zellers et al. 2019)).
However, as LLM performance continues to improve, several mainstream benchmarks have approached saturation, with multiple LLMs achieving near-perfect scores. This diminishes the discriminative power and evaluation value of these benchmarks, forcing the community to develop new benchmarks frequently. Meanwhile, since test datasets have been publicly available and fixed for a long time, white-box evaluation is also facing an increasingly severe issue of data contamination. That is, LLMs may have encountered the test data during training, thereby compromising the credibility of evaluation results (Yang et al. 2023; Kiela et al. 2021). Several studies have raised concerns about this issue (Li et al. 2024). For example, the Apple team (Mirzadeh et al. 2024) found that even minor alterations to mathematical question values or the insertion of irrelevant information can significantly reduce LLM accuracy. Similarly, Xie (Xie et al. 2024) changed the underlying principles of questions while maintaining semantic similarity. They observed that LLMs still tended to apply the original solutions, exhibiting “memorization” rather than reasoning behavior.
Black-Box
To alleviate the issue of data contamination, researchers have proposed black-box evaluation. This approach involves constructing entirely new test datasets and keeping these datasets undisclosed, or releasing them only after the evaluation phase, thus reducing the risk of the data being incorporated into training data. For example, Huang (Huang et al. 2023) evaluated LLMs using real-time data from weekly Codeforces competitions. Big-Bench (Srivastava et al. 2023) only releases a subset of its tasks before evaluation to prevent data leakage. GPQA (Rein et al. 2024) utilizes a newly created question bank manually composed by human experts. As a typical representative of black-box evaluation, SuperGLUE (Wang et al. 2019) requires LLMs to be uploaded to the platform, where evaluation is conducted in a completely closed environment, and only the evaluation results are returned. This fundamentally prevents the test set from being included in the LLM’s training data.
Although black-box evaluation enhances evaluation confidentiality, its inherent black-box operation feature also leads to issues of insufficient verifiability. Since the test data and evaluation process are not visible to researchers, the objectivity and professionalism of the evaluation are difficult to independently verify, resulting in a lack of credibility and interpretability. Research (Casper et al. 2024) has explicitly pointed out that relying solely on LLM outputs is insufficient to meet rigorous evaluation requirements. Furthermore, datasets used in black-box evaluation may still originate from publicly available internet data or may have been leaked and incorporated into the training data before evaluation. In other words, as long as the datasets remain fixed, the risk of data contamination persists.
Human-Crowdsourced
To fundamentally avoid the data contamination caused by fixed test datasets, researchers from the University of California, Berkeley, proposed a human preference-based crowdsourced platform called Chatbot Arena (Chiang et al. 2024). This platform integrates multiple LLMs, where different LLMs independently answer user questions. Ultimately, users compare the responses and indicate their subjective preference, thereby generating a ranking.
Although this method possesses inherent dynamism, it still has some drawbacks. First, the evaluation relies heavily on users’ subjective preference, which are prone to bias (Blunch 1984; Raghubir and Valenzuela 2006; Tuch et al. 2012; Hamborg, Hülsmann, and Kaspar 2014). In addition, research (Connor Chen and Tianle Li 2025; Li, Angelopoulos, and Chiang 2024) has found that responses that are longer, contain emojis, or are organized as points tend to be favored. Second, compared to professionally designed test datasets by experts, user-generated questions in human-crowdsourced evaluation vary widely in quality. It is also unrealistic to require every user to possess domain expertise during the evaluation process. Last but not least, although Chatbot Arena discloses its ranking rules and metric computation methods, its core raw data (such as user inputs, LLM outputs, and evaluation records) is not made public. As a result, key aspects of the evaluation process remain opaque, making it difficult to verify or review independently.
Figure 2: Evaluation pipeline overview.
Evaluation Criteria
In summary, we believe that a scientific evaluation of LLMs should satisfy four key evaluation criteria.
1. Dynamic: Each evaluation should use questions and answers that possess adequate randomness (freshness) to mitigate the issue of data contamination.
2. Transparent: The evaluation methods should be open and transparent, and the evaluation process and results should be reproducible by anyone, thereby increasing public credibility.
3. Objective: The evaluation standards and results should not be influenced by individual subjective preference, thus reducing subjective bias and the possibility of manipulation, to accurately reflect the objective performance of the LLMs.
4. Professional: The questions and answers employed in evaluations should be at or near the average level of human experts in relevant domains, enabling in-depth evaluation of the LLM’s professional capabilities in vertical domains.
None of the existing evaluation methods can simultaneously satisfy these four criteria. Therefore, we propose LLM-Crowdsourced.
LLM-Crowdsourced
According to Figure 1, white-box evaluation (based on public datasets) suffers from a lack of dynamism, black-box evaluation (based on unknown datasets) is deficient in transparency, and human-crowdsourced evaluation (based on human-crowdsourced) is subject to subjective preference, incomplete transparency, and insufficient professionalism. To address these issues, we propose an evaluation paradigm based on LLM-Crowdsourced. In this paradigm, multiple LLMs under evaluation generate question, answer independently, and evaluate mutually, with final capability rankings computed by aggregating scores across rounds.
Evaluation Pipeline
As illustrated in Figure 2, our method can be divided into four phases. Before the evaluation begins, it is necessary to predefine the evaluation domain and clarify the corresponding rules to ensure controllability and consistency in both question generation and evaluation processes.
Phase 1 Generate Question: Each LLM takes turns acting as the questioner, generating an original and difficult question according to the predefined rules, and providing a reference answer. Subsequently, this question is distributed to the other LLMs for answering.
Phase 2 Answer Independently: Except for the questioner, the remaining LLMs independently answer the question without referencing others’ answers, thereby evaluating their capabilities.
Phase 3 Evaluate Mutually: The question-generating LLM provides the reference answer for the other LLMs. Based on the reference answer and the scoring criteria, each LLM evaluates the answers of other LLMs (excluding its own), thus enabling mutual evaluation among the LLMs.
Phase 4 Update Ranking: After aggregating the mutual evaluation results for the current round, each LLM’s score for the round is calculated, and the ranking is updated in real-time. This stage does not represent the final ranking; rather, it serves as periodic feedback and provides a basis for evaluations in subsequent rounds. After n rounds, the final ranking for all LLMs is output.
Advantages
The proposed LLM-Crowdsourced evaluation paradigm aims to integrate the strengths of existing evaluation methods while overcoming their inherent limitations, thus establishing a more scientific and standardized evaluation system. This approach aligns with four key evaluation criteria for scientific LLM evaluation as discussed in the Evaluation Criteria section.
In terms of dynamism, both the evaluation questions and scoring criteria are generated by LLMs based on established rules. This ensures that each evaluation is entirely new content, thereby avoiding overlap with training data. This mechanism effectively mitigates issues such as benchmark saturation over time and data contamination, enhancing the long-term validity of the evaluation.
In terms of transparency, question generation, answering procedures, and scoring criteria are all made public, with every step traceable and easily reproducible. This enables external researchers to independently verify the evaluation process and results, thereby increasing the interpretability and credibility of the evaluation system.
In terms of objectivity, the approach employs a decentralized mutual evaluation mechanism, whereby multiple LLMs jointly perform the evaluation task. This reduces the influence of individual LLM preferences or human subjective preference (Chen et al. 2024), enhancing the fairness and stability of the evaluation outcomes.
In terms of professionalism, LLMs, compared to human-crowdsourced evaluation, possess a more professional background knowledge, enabling them to deliver more specialized and reliable results in evaluation tasks.
In summary, our evaluation approach not only addresses the current challenges faced by LLM evaluation, such as high subjectivity, difficulties in review, and lack of generalizability, but also provides new ideas and references for establishing unified evaluation standards in the future.
Principle of Experiment Design
In this paper, we propose a general evaluation paradigm for LLMs that allows for flexible configuration of evaluation procedures and rules tailored to the specific requirements of individual domains, thereby accommodating the diverse capability needs of different tasks. To validate the feasibility and adaptability of this paradigm, we selected two representative capability dimensions, mathematics and programming, and conducted three sets of experiments, respectively. Detailed descriptions are provided in subsequent sections. Our selection of these two domains is primarily motivated by two considerations: (1) These capabilities are currently among the most essential and frequently utilized core features by LLM users; (2) These domains align with fundamental disciplines in general human education, making them broadly representative and generalizable. In the experimental setup, we devised specific question generation rules and scoring mechanisms for each domain within a unified evaluation framework.
In this study, we employed the unified API provided by Huiyan Corporation to invoke the official APIs of multiple LLMs. We selected eight of the most advanced LLMs currently available (with n=8 in subsequent experiments). The LLMs and their respective version numbers are as follows: gpt-4.1-2025-04-14, o3-mini, o1, claude-3-7-sonnet-20250219, deepseek-r1, deepseek-v3-0324, qwen2.5-max, and gemini-2.5-pro-exp-03-25.
Figure 3: Prompts design using role-play and zero-shot for professional and dynamic mathematics question generation.
LLM-Crowdsourced On Mathematics
Experiment Design
Question Generation Mechanism Design
Mathematics, as a highly formalized language, possesses rigorous reasoning rules and objective standard answers, making it inherently suitable for evaluating the logical reasoning, abstract modeling, and generalization capabilities of LLMs. To mitigate the issue of data contamination caused by static question banks, we have designed a more professional and dynamic mathematical evaluation framework. In the question generation phase (see Figure 3), the question-generating LLM is tasked with acting as a mathematician and formulating professional questions in the mathematical domain. Recognizing that LLMs have encountered substantial mathematical content during training (Achiam et al. 2023), we do not restrict the question type. We require the LLM to identify the most difficult questions based on its prior knowledge, abstract the core principles, and construct an original question accordingly, while also providing a complete reference solution. This mechanism not only increases the novelty and difficulty of the questions but also accentuates the capability differences among LLMs and differentiates between “memorization-based answering” and “reasoning-based answering.”
Evaluation Mechanism Design
Since mathematical questions typically allow for various reasonable solutions, traditional scoring methods struggle to capture different lines of reasoning and are prone to introducing subjective bias from the LLMs. To address this, we employ a relative evaluation mechanism. That is, each LLM evaluates others based on the reference solutions, and the resulting rankings are normalized into final scores. This approach mitigates the subjective bias introduced by absolute scoring standards and promotes more stable and equitable comparisons in multi-LLM evaluations.
Experiment Results
Figure 4 shows that the eight mainstream LLMs exhibit significant differences in their performance on mathematical tasks. The figure presents the scores and fluctuation ranges for each LLM in multi-turn question answering, reflecting overall performance and answering stability. It can be observed that LLMs with leading scores, such as gemini-2.5-pro-exp and gpt-4.1, demonstrate balanced performance across various questions, exhibiting strong and stable logical reasoning abilities. In contrast, lower-ranked LLMs, such as o3-mini and o1, show large score fluctuations and evident weaknesses. Overall, the figure intuitively displays the differences in mathematical capabilities among the LLMs and also verifies the effectiveness of this evaluation paradigm in capability differentiation.
Figure 4: Radar charts of LLMs’ performance across multiple mathematics questions.
Findings 1: Some LLMs demonstrate highly professional mathematical question-design capabilities. There are pronounced differences among the LLMs in their ability to generate questions. Some LLMs, such as gemini-2.5-pro-exp, demonstrated professional-level mathematical creativity, as illustrated by the following question:
Let $S_{3}(n)$ denote the sum of the digits of a positive integer $n$ when written in base 333. Define a sequence $a_{n}=(-1)^{S_{3}(n)}$, for $n\geq 1$.Consider the Dirichlet series $F(s)=\sum_{n=1}^{\infty}\frac{a_{n}}{n^{s}}$ , defined for complex numbers $s$ where the series converges. Determine if $F(s)$ can be analytically continued to a region including $s=0$, and if so, find the value $F(0)$.
This question merges elements from number systems, automatic sequences, and the analytic continuation of complex analysis, showcasing a high degree of innovation and theoretical challenge. Its design approach is highly consistent with the method for constructing Dirichlet series for the Riemann zeta function in modern number theory. It constructs an original question that is difficult to find in public datasets or human-crowdsourced evaluations. Although most LLMs were unable to provide the correct solution, the differentiation in their performance substantiates the effectiveness of the “LLM-prior-knowledge-based question-posing mechanism” described in the Experiment Design Section, which enables distinguishing capability differences among LLMs and evaluating their generalization abilities. This also highlights the advantages of our method in terms of professionalism and dynamism.
Findings 2: “Memorization-based answering” behavior does indeed exist in LLMs. In experiments, we observed the phenomenon of “memorization-based answering,” where the LLM incorrectly treats a question as one with a similar structure but fundamentally different content. For instance, in the Dirichlet series question mentioned above, deepseek-r1 attempted to formalize $F(s)$ as an infinite product structure similar to an Euler product, and thus derived an incorrect result. Euler products commonly use this approach for handling Dirichlet series such as the Riemann zeta function. However, since the problem’s assumptions in this case do not satisfy the requirements of a multiplicative structure, this solution is not applicable. This mistake reveals the typical characteristics of “memorization-based answering” behavior. Similar mistakes appeared 10 times among the 168 responses in our mathematics experiments, further confirming the limitations of white-box evaluation methods; LLMs may match questions based on formal similarity, lacking true reasoning ability. In contrast, our method can effectively identify such “memorization-based answering” behaviors and reveal the true boundaries of the LLM’s reasoning capabilities.
Findings 3: LLM evaluation results demonstrate high consistency (robustness). During the evaluation phase, we found that multiple LLMs exhibited high consistency in evaluating the answers of other LLMs. For example, GPT-4.1’s answers were consistently rated by several LLMs as having the clearest logical chain, and the answers from gemini-2.5-pro-exp were frequently deemed superior to the reference answers. Top-k consistency analysis shows that when $k=n/2$, the Top-k value is 74.85%, indicating that most LLMs maintain stable evaluation standards when evaluating other LLMs, thus forming a cross-LLM “consensus judgment standard.” This result also indirectly confirms the robustness and objectivity of our method.
LLM-Crowdsourced On Programming
Figure 5: Prompts design using role-play and few-shot for professional and dynamic programming question generation.
Experiment Design
Question Generation Mechanism Design
Programming tasks involve both theoretical complexity and engineering implementation, yet existing benchmarks primarily focus on code output accuracy, seldom addressing question design, which makes it challenging to reflect a LLM’s capability in program synthesis comprehensively. To address this, we propose a professional and dynamic programming evaluation framework. As shown in Figure 5, in each round, the question-generating LLM acts as a senior programming competition question setter, designing professional-level questions within the programming domain. By studying high-level competition questions (such as ACM), the LLM masters structural design and complexity factors of high-quality programming questions. Following this style, the LLM generates a brand-new, non-derivative, and difficult original programming question, and provides the question description, input/output format, constraint specification, reference solution, and complexity analysis. During the answering phase, LLMs can freely choose the programming language, allowing them to leverage their strengths while eliminating biases from artificial constraints.
Evaluation Mechanism Design
To better reflect professionalism and objectivity, all LLMs are required to evaluate the answers of others based on the reference code and evaluation criteria. The evaluation mechanism employs a 100-point scale, which comprehensively considers multiple dimensions, including correctness, execution efficiency, readability, structural clarity, and memory control. The LLMs conduct cross-evaluation, further enhancing the professional interpretability and objectivity of the evaluation results.
Figure 6: Dual-axis comparison of LLMs on programming tasks: solving ability vs. question difficulty.
Experiment Results
Findings 4: The consistency between LLMs’ question-setting and question-solving abilities. As shown in Figure 6, we conducted a two-dimensional analysis of eight LLMs’ performance in programming tasks: the average score a LLM receives on other LLMs’ questions (Own Average Score on Others’ Questions) measures its question-solving ability, while the average score other LLMs receive on that LLM’s questions (Others’ Average Score on Own Question) reflects the difficulty of the questions constructed by the LLM. The overall results reveal significant performance disparities among LLMs in both question-setting and question-solving phases, and LLMs with strong comprehensive abilities tend to perform well in both aspects.
Taking gemini-2.5-pro-exp-03-25 as an example, the questions it generates often involve cutting-edge topics with considerable theoretical depth, such as quantum path counting, significantly raising the threshold and discriminative power of the questions. In the answering phase, gemini demonstrated capabilities in formulaic derivation and complexity optimization on multiple occasions. For example, in the “Cube Garden” question, while most other LLMs employed brute-force approaches (with a time complexity of $O(K^{3})$), gemini derived a closed-form solution via algebraic induction as $K^{3}-3K+2$, efficiently optimizing the complexity to $O(1)$. In contrast, LLMs such as qwen2.5-max proposed questions that other LLMs can easily achieve high scores, reflecting insufficient challenge or discriminative power. Additionally, these LLMs often rely on brute-force enumeration. They were prone to omissions in handling boundary conditions and abnormal inputs, exhibiting deficiencies in algorithmic generalization and engineering implementation capabilities.
Discussion and Limitations
Q1: Do the LLMs generate the questions and answers reliable?
We require question-generating LLMs to provide reference answers simultaneously, ensuring each question has a clear, verifiable solution. All questions underwent manual review, with no unsolvable or logically flawed items found. Even if reference answers have flaws, LLMs can independently evaluate answer quality, avoiding reliance on a single reference. With multiple LLMs rotating as questioners, even low-discrimination questions can be diluted through cross-evaluation. The consistent performance of high-scoring LLMs and the Top-k consistency analysis both confirm the stability of the evaluation.
Q2: If the evaluation LLMs may have biases or errors, how can the results be ensured to be objective and reasonable?
We prohibit self-evaluation and adopt a “cross-evaluation + reference answer” mechanism to reduce bias. Even if misjudgments occur, scores are corrected through multiple rounds of cross-evaluation. We perform three independent runs per experiment to mitigate LLM-induced noise, and the scoring results have demonstrated high consistency and robustness. Among the eight selected LLMs, even the weakest outperforms human-crowdsourced on average. Compared with mainstream benchmarks and Chatbot Arena, although there are slight differences in orderings, our results show consistent overall trends, further confirming reliability. Although evaluation LLMs are not perfect, our goal is to explore a scalable, human-free evaluation paradigm.
Conclusion
To overcome the challenges posed by data contamination, black-box operation, and subjective preference present in mainstream evaluation paradigms, such as white-box evaluation, black-box evaluation, and human-crowdsourced evaluation, this paper proposes LLM-Crowdsourced, a novel benchmark-free evaluation paradigm. This approach possesses four core advantages: dynamic, professional, objective, and transparent. Experiments on two classical domains, mathematics and programming, demonstrate that LLM-Crowdsourced can not only effectively distinguish differences in LLM capability, but also identify phenomena that are difficult to detect using traditional methods, such as “memorization-based answering.” The approach exhibits excellent discriminability and generalizability. We have publicly released the complete evaluation framework and experimental code to support community evaluation of new LLMs and to promote the standardization and advancement of LLM evaluation methods.

### v2

@@
 Experiment Design
-Question Generation Mechanism Design
-Mathematics, as a highly formalized language, possesses rigorous reasoning rules and objective standard answers, making it inherently suitable for evaluating the logical reasoning, abstract modeling, and generalization capabilities of LLMs.
+Question Generation Mechanism Design: Mathematics, as a highly formalized language, possesses rigorous reasoning rules and objective standard answers, making it inherently suitable for evaluating the logical reasoning, abstract modeling, and generalization capabilities of LLMs.
 To mitigate the issue of data contamination caused by static question banks, we have designed a more professional and dynamic mathematical evaluation framework.
@@
 This mechanism not only increases the novelty and difficulty of the questions but also accentuates the capability differences among LLMs and differentiates between “memorization-based answering” and “reasoning-based answering.”
-Evaluation Mechanism Design
-Since mathematical questions typically allow for various reasonable solutions, traditional scoring methods struggle to capture different lines of reasoning and are prone to introducing subjective bias from the LLMs.
+Evaluation Mechanism Design: Since mathematical questions typically allow for various reasonable solutions, traditional scoring methods struggle to capture different lines of reasoning and are prone to introducing subjective bias from the LLMs.
 To address this, we employ a relative evaluation mechanism.
-That is, each LLM evaluates others based on the reference solutions, and the resulting rankings are normalized into final scores.
+Each LLM evaluates others based on the reference solutions, and the rankings are normalized into final scores.
 This approach mitigates the subjective bias introduced by absolute scoring standards and promotes more stable and equitable comparisons in multi-LLM evaluations.
@@
 Figure 4 shows that the eight mainstream LLMs exhibit significant differences in their performance on mathematical tasks.
-The figure presents the scores and fluctuation ranges for each LLM in multi-turn question answering, reflecting overall performance and answering stability.
+The figure presents the scores and fluctuations for each LLM in multi-turn question answering, reflecting overall performance and answering stability.
 It can be observed that LLMs with leading scores, such as gemini-2.5-pro-exp and gpt-4.1, demonstrate balanced performance across various questions, exhibiting strong and stable logical reasoning abilities.
 In contrast, lower-ranked LLMs, such as o3-mini and o1, show large score fluctuations and evident weaknesses.
-Overall, the figure intuitively displays the differences in mathematical capabilities among the LLMs and also verifies the effectiveness of this evaluation paradigm in capability differentiation.
+Overall, the figure intuitively displays the differences in mathematical capabilities among the LLMs and verifies the effectiveness of our method in capability differentiation.
 Figure 4: Radar charts of LLMs’ performance across multiple mathematics questions.
@@
 For example, gpt-4.1’s answers were consistently rated by several LLMs as having the clearest logical chain, and the answers from gemini-2.5-pro-exp were frequently deemed superior to the reference answers.
-Top-k consistency analysis shows that when $k=n/2$, the Top-k value is 74.85%, indicating that most LLMs maintain stable evaluation standards when evaluating other LLMs, thus forming a cross-LLM “consensus judgment standard.” This result also indirectly confirms the robustness and objectivity of our method.
+Top-k consistency analysis shows that when $k=n/2$, the Top-k value is 74.85%, indicating that most LLMs maintain stable evaluation standards when evaluating other LLMs, thus forming a cross-LLM “consensus judgment standard.” Moreover, the standard deviations of the scores remain low, ranging from 2.04 to 9.87 on a 100-point scale.
+These results quantitatively demonstrate the robustness and stability of the evaluation results, even with a limited sample size.
 LLM-Crowdsourced On Programming
@@
 Experiment Design
-Question Generation Mechanism Design
-Programming tasks involve both theoretical complexity and engineering implementation, yet existing benchmarks primarily focus on code output accuracy, seldom addressing question design, which makes it challenging to reflect a LLM’s capability in program synthesis comprehensively.
+Question Generation Mechanism Design: Programming tasks involve both theoretical complexity and engineering implementation, yet existing benchmarks primarily focus on code output accuracy, seldom addressing question design, which makes it challenging to reflect a LLM’s capability in program synthesis comprehensively.
 To address this, we propose a professional and dynamic programming evaluation framework.
@@
 During the answering phase, LLMs can freely choose the programming language, allowing them to leverage their strengths while eliminating biases from artificial constraints.
-Evaluation Mechanism Design
-To better reflect professionalism and objectivity, all LLMs are required to evaluate the answers of others based on the reference code and evaluation criteria.
+Evaluation Mechanism Design: To better reflect professionalism and objectivity, all LLMs are required to evaluate the answers of others based on the reference code and evaluation criteria.
 The evaluation mechanism employs a 100-point scale, which comprehensively considers multiple dimensions, including correctness, execution efficiency, readability, structural clarity, and memory control.
@@
 We require question-generating LLMs to provide reference answers simultaneously, ensuring each question has a clear, verifiable solution.
-All questions underwent manual review, with no unsolvable or logically flawed items found.
+All questions and references were manually reviewed by at least three humans, with no unsolvable or logically flawed items found.
 Even if reference answers have flaws, LLMs can independently evaluate answer quality, avoiding reliance on a single reference.
 With multiple LLMs rotating as questioners, even low-discrimination questions can be diluted through cross-evaluation.
-The consistent performance of high-scoring LLMs and the Top-k consistency analysis both confirm the stability of the evaluation.
+The high Top-k consistency and the low standard deviations both quantitatively confirm the robustness and stability of the evaluation results, even with a limited sample size.
 Q2: If the evaluation LLMs may have biases or errors, how can the results be ensured to be objective and reasonable?
@@
 We perform three independent runs per experiment to mitigate LLM-induced noise, and the scoring results have demonstrated high consistency and robustness.
-Among the eight selected LLMs, even the weakest outperforms human-crowdsourced on average.
+We do not claim to offer a perfect solution.
+Instead, we propose a scalable, human-free evaluation paradigm to explore more objective and reproducible evaluation methods.
+Despite the exploratory nature and limited sample size, our results already exhibit strong robustness and lay the foundation for future large-scale studies.
+Among the eight selected LLMs, even the weakest outperforms humans on average.
 Compared with mainstream benchmarks and Chatbot Arena, although there are slight differences in orderings, our results show consistent overall trends, further confirming reliability.
-Although evaluation LLMs are not perfect, our goal is to explore a scalable, human-free evaluation paradigm.
 Conclusion
 To overcome the challenges posed by data contamination, black-box operation, and subjective preference present in mainstream evaluation paradigms, such as white-box evaluation, black-box evaluation, and human-crowdsourced evaluation, this paper proposes LLM-Crowdsourced, a novel benchmark-free evaluation paradigm.
-This approach possesses four core advantages: dynamic, professional, objective, and transparent.
-Experiments on two classical domains, mathematics and programming, demonstrate that LLM-Crowdsourced can not only effectively distinguish differences in LLM capability, but also identify phenomena that are difficult to detect using traditional methods, such as “memorization-based answering.” The approach exhibits excellent discriminability and generalizability.
-We have publicly released the complete evaluation framework and experimental code to support community evaluation of new LLMs and to promote the standardization and advancement of LLM evaluation methods.
+It integrates four key evaluation criteria: dynamic, transparent, objective, and professional.
+Experiments on two classical domains, mathematics and programming, demonstrate that LLM-Crowdsourced can not only effectively distinguish differences in LLM capability, but also identify phenomena that are difficult to detect using traditional methods, such as “memorization-based answering.” This method exhibits excellent discriminability and generalizability.
+We have publicly released the complete evaluation framework and code to support community evaluation of new LLMs and to promote the standardization and advancement of LLM evaluation methods.